OpenStudy (anonymous):

WHICH OF THE FOLLOWING ARE EQUAL TO THE EXPRESSION BELOW ? CHECK ALL THAT APPLY SQRT 44/2 A.SQRT 44/4 B.SQRT11 C.22 D.SQRT22???

OpenStudy (anonymous):

B

OpenStudy (anonymous):

\[\sqrt{\frac{44}{2}}=\sqrt{22} \]

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Yeah it's D. I calculated wrong. My bad!